# Computing Anharmonic Infrared Spectra of Polycyclic Aromatic Hydrocarbons Using Machine-Learning Molecular Dynamics

### v1

@@ Abstract
Polycyclic aromatic hydrocarbons (PAHs) are key contributors to interstellar infrared emission bands (AIBs). However, current spectral databases for AIB analysis face limitations because of the neglect of anharmonicity in vibrations and of temperature effects, which result from the high computational cost of traditional quantum chemical calculations (QCCs). In this work, we present a machine-learning-based molecular dynamics (MLMD) approach that efficiently computes anharmonic infrared (IR) spectra while accounting for temperature effects. MLMD achieves predictive accuracy comparable to that of QCCs but with significantly reduced computational cost, scaling linearly with the number of atoms in the system. We apply MLMD to calculate the anharmonic spectra of 1 704 PAHs in the NASA Ames PAH IR Spectroscopic Database with up to 216 carbon atoms, demonstrating its ability to handle large molecular systems. Our results suggest that MLMD can facilitate the creation of extensive molecular spectral databases, advancing AI-assisted analyses of astronomical IR spectra, particularly with upcoming data from the James Webb Space Telescope.
@@ I Introduction
In the mid-1970s, distinct infrared (IR) signals were identified in the universe, notably at wavelengths of 3.3, 6.2, 7.7, 8.6, 11.2, and 12.7 $\upmu$m. These signals were later linked to the vibrations of C-H and C-C bonds in aromatic molecules, known as aromatic IR bands (AIBs) (Leger & Puget, 1984; Allamandola et al., 1985). Such emissions are primarily attributed to polycyclic aromatic hydrocarbons (PAHs) after the absorption of ultraviolet (UV) photons. PAHs are recognized as key contributors to the evolution of the interstellar medium, influencing processes such as gas heating, ionization balance, and star formation (Tielens, 2008). Determining the population, composition, size distribution, charge state, and chemical structures of interstellar PAHs provides valuable insight into the physical conditions of their host galaxies. However, this remains a significant challenge because of the structural diversity of PAHs. Variations in carbon-ring arrangements, side groups, substitutions, charge states, and isotopologues lead to distinct IR emission characteristics, each governed by specific vibration modes (Peeters et al., 2021).
In order to tackle these challenges, quantum chemical calculations (QCCs) are now an indispensable aid in AIBs analysis. Using QCCs, publicly available spectral databases have been established for AIB analysis, such as the NASA Ames PAH IR Spectroscopic Database (PAHdb) (Mattioda et al., 2020; Bauschlicher et al., 2018; Boersma et al., 2014). The availability of these databases enables extensive data-driven analysis of AIBs, providing a more robust interpretation of their features (Boersma et al., 2013, 2015; Maragkoudakis et al., 2020; Sadjadi et al., 2015; Cami, 2011; Shannon & Boersma, 2019; Ricca et al., 2021). Furthermore, these databases have facilitated the analysis of astrochemical components using machine learning (ML) models, demonstrating significant computational efficiency (Kovács et al., 2020) and helping to uncover relationships between molecular structures and spectra (Meng et al., 2021, 2023). Despite these advancements, current theoretical spectral databases face three primary limitations:
@@
  • Neglect of temperature effect: AIB features are typically attributed to IR emission from far-UV-pumped PAHs, where an absorbed UV photon can elevate the effective molecular temperature to several hundred or even thousands of kelvin, an effect that is not taken into account in most theoretical spectral data.
@@
These challenges arise mainly from the high computational cost of QCCs. In particular, the computation time for anharmonic molecular spectra using second-order vibrational perturbation theory (VPT2) is prohibitively long, typically only accommodating molecules with $N_{\text{C}}<25$, where $N_{\text{C}}$ represents the number of carbon atoms in the system (Mackie et al., 2016; Esposito et al., 2024a). The cost of incorporating temperature effects into these calculations using the Wang-Landau random walk algorithm is even more prohibitive, being orders of magnitude higher due to the need for extensive sampling (Chen, 2018; Chen et al., 2019). Facing the vast diversity of PAH species, these limitations significantly hinder the accurate interpretation of observational data, posing substantial challenges in the era of advanced observational facilities such as the James Webb Space Telescope (Boersma et al., 2023; Ricca et al., 2024).
@@ II.1 Training Data
The chemical structures of 687 neutral PAH molecules were extracted from the theoretical dataset of PAHdb (v3.2, Mattioda et al. (2020)) to construct the training datasets. The selected PAHs encompass a diverse array of chemical structures that are of astronomical significance. Classical MD simulations were performed using a custom code to generate random configurations of the atoms in each molecule at 300 K, using the adaptive interatomic reactive empirical bond order potential (Stuart et al., 2000). In these simulations, the molecules reached thermal equilibrium in 200 ps using a canonical Nose-Hoover thermostat, with a time step of 0.5 fs. Ten extended structures were extracted from the atomic trajectories of each molecule. As these geometries do not necessarily fall near equilibrium configurations, random Gaussian perturbations were applied to the ground-state atomic coordinates to create 15 more configurations for each molecule to improve the generalizability of the model. Finally, 17 175 configurations (also called molecular conformations) were generated in total.
@@ II.2 NNFF
The NNFF utilizes a deep residual network architecture (ResNet) (He et al., 2016a), implemented within the JAX framework (Bradbury et al., 2018). A ResNet employs skip connections to learn the residual function between the input and output layers, effectively addressing the data degradation problem that arises with increased network depth (He et al., 2016b). This feature is particularly beneficial for modeling complex molecules like large PAHs, as it enables the capture of deep dependencies and interaction patterns among atoms through multilevel learning, which ultimately leads to more accurate predictions of physicochemical properties (Zhang et al., 2024; Xue et al., 2024).
@@
$\begin{split}L&=0.99\left\langle\frac{0.2}{N_{\text{atom}}}\sum^{N_{\text{atom}}}\mathrm{ln}\left[\mathrm{cosh}\left(\frac{\left\|\mathbf{f}_{\text{pred}}-\mathbf{f}_{\text{ref}}\right\|_{2}}{0.2}\right)\right]\right\rangle\\ &+0.01\left\langle 0.02\ \mathrm{ln}\left[\mathrm{cosh}\left(\frac{E_{\text{pred}}-E_{\text{ref}}}{N_{\text{atom}}\times 0.02}\right)\right]\right\rangle,\end{split}$ (1)
where ${N_{\text{atom}}}$ represents the total number of atoms in a molecule. The subscripts ${}_{\text{pred}}$ and ${}_{\text{ref}}$ denote the values predicted by NNFF and DFT, respectively. $\left\|\mathbf{f}_{\text{pred}}-\mathbf{f}_{\text{ref}}\right\|_{2}$ represents the Euclidean norm of the vector difference between the predicted force and the reference force. The model adopts a core-width sequence of 64:32:16:16 and employs the fully non-linear VeLO optimization technique (Metz et al., 2022) for 600 epochs of training to enhance convergence (Carrete et al., 2023).
To evaluate the performance of the model, we conducted tests on a set of configurations for 3 580 PAHs that were excluded from the training set. The results presented in Figure 2b demonstrate that the NNFF model accurately reproduces the forces calculated by DFT, achieving a coefficient of determination (${R^{2}}$) of approximately 0.99 and a mean RMSE of 0.00117 hartree/bohr (see Table 1), indicating strong generalizability to unseen samples. This capability allows the NNFF model to effectively replace DFT electronic structure calculations in AIMD, providing robust support for subsequent IR spectrum calculations.
@@ II.3 EPNN
where $p^{\left(\alpha\right)}_{i,\text{pred}}$ and $p^{\left(\alpha\right)}_{i,\text{ref}}$ represent the predicted and reference values of the $\alpha$-th Cartesian component of $\mathbf{p}$ for the $i$-th molecule. $N_{\text{mol}}$ denotes the total number of molecules. To evaluate the performance of the model, we predicted the $\mathbf{p}_{\text{ref}}$ of 3 580 PAH conformations excluded from the training set. As shown in Figure 2c, there is strong agreement between the predicted and reference values, with a ${R^{2}}$ reaching 0.97 and a mean RMSE of 0.0545 debye, as shown in Table 1.
@@ II.4 Anharmonic IR spectrum
To compute the anharmonic IR spectrum of a PAH, MD simulations were performed to simulate its finite-temperature dynamics based on atomic forces predicted by the trained NNFF. We used the MD implementation in the Atomic Simulation Environment (ASE) (Larsen et al., 2017). Initially, the molecule reaches thermal equilibrium in 200 ps in a canonical ensemble (NVT) at a target temperature controlled by the Nosé-Hoover thermostat, with a timestep of 0.5 fs. The system was then further simulated in a microcanonical ensemble (NVE) for 200 ps, recording atomic configurations at every 1.0 fs during vibration.
These configurations serve as input to the EPNN, which is used to compute $\mathbf{p}(t)$ at each time step of the vibration. The dipole time autocorrelation function, characterizing the time evolution of $\mathbf{p}$, is given by:
@@
where $\tau$ denotes the total simulation time.
The autocorrelation function $\langle\mathbf{p}(0)\cdot\mathbf{p}(t)\rangle$ is then subjected to a Fourier transform to obtain the IR intensity:
[…]
where $\omega$ denotes the angular frequency, $k_{\text{B}}$ represents the Boltzmann constant, $\hbar$ is the reduced Planck constant, $T$ indicates the temperature, and $t$ is time.
@@ II.5 Reference methods for comparison
To evaluate the accuracy of our model, we compared it with two established QCC methods to calculate the IR spectra of PAHs. The first method, designated as DFT harmonic, is a widely adopted hybrid DFT approach for harmonic spectrum calculations. In this method, the vibrational frequencies (normal modes) are calculated from the second derivatives of the potential energy with respect to the positions of the nuclei at the stationary point of the system. IR intensities are determined using the double harmonic approximation, which involves computing the derivatives of $\mathbf{p}$ with respect to the normal modes (Langhoff, 1996). The calculations were carried out at the B3LYP/4-31G level of theory using Gaussian 16, ensuring consistency with the QCCs employed in our training dataset. It is important to note that the assumption of a harmonic potential frequently underestimates fundamental frequencies, often necessitating empirical scaling factors to align the computational results with the experimental data (Bauschlicher & Langhoff, 1997). Despite the transferability issues introduced by these empirical scaling factors, this method remains the most commonly utilized method for molecular IR spectrum calculations due to its relatively low computational cost. For the sake of a fair comparison, we do not apply any scaling factors to the DFT harmonic results, as these factors should ideally be derived from fitting to experimental spectra. The DFT harmonic method was employed to compute the harmonic spectra for 49 PAH molecules, for which experimentally measured spectra are available in the PAHdb.
@@
The second method used for comparison is the VPT2 anharmonic spectrum calculations combined with B3LYP/6-311+g* DFT, referred to here as DFT anharmonic. This approach represents the state-of-the-art for computing the molecular anharmonic IR spectra, and has been implemented in Gaussian 16. It requires the computation of the second, third, and fourth derivatives of the potential, which results in significant computational demands (Nielsen, 1951; Mackie et al., 2015). Due to this cost, we applied VPT2 to calculate the anharmonic spectra of 26 PAHs, with the maximum $N_{C}$ being 32, from the 49 PAHs used for the DFT harmonic method. In the comparison between the computed and experimental spectra, the set of discrete lines in the IR spectrum is turned into a histogram of lines within bins with a width of $16.65\text{cm}^{-1}$ determined using Knuth’s Bayesian rule (Knuth, 2006).
@@
The 49 PAHs for which the experimental spectra are available in PAHdb (v3.2), ranging from 10 to 50 carbon atoms, exhibit a variety of chemical structures, as detailed in Appendix I. The experimental IR spectra were obtained using matrix isolation techniques at low temperatures (Hudgins & Allamandola, 1995; Hudgins & Sandford, 1998a, b, c; Hudgins & Allamandola, 1999; Mattioda et al., 2003, 2005, 2014, 2017). To facilitate comparison with experimental data, we employed theoretical spectra computed using MLMD at an effective temperature of 50 K, which accounts for molecular vibrations under experimental conditions (Esposito et al., 2024b).
@@ III Results and discussion
Figure 3: Comparison of IR spectra for five PAHs computed using the MLMD at 50 K(left, red), DFT harmonic (middle, green), and DFT anharmonic (right, blue) methods, alongside the experimentally measured spectra (black). The intensities are normalized to the maximum values across the four spectra for ease of comparison. The root mean square error (RMSE) is computed from the original intensity values, with units of $10^{5}\text{cm}^{2}/\text{mol}$.
@@
This dataset is expected to be a valuable resource for the spectral decomposition of AIBs and for data-driven studies investigating the precise structure-spectral relationships of PAHs. In addition, it will aid in the exploration of temperature effects. For instance, Figure 7 displays the anharmonic IR spectra of a simple 0mixture of the 1 704 neutral PAHs at 50 K (blue), 300 K (green), and 600 K (red). These spectra represent the sum of individual PAH spectra and are compared to the ground-state harmonic spectra from PAHdb (gray). Notably, the IR signals are significantly enhanced at high temperatures, exhibiting stronger peaks and an elevated plateau. A redshift in peak positions is also observed at high temperatures as a general trend. Furthermore, some emission bands that are weak at lower temperatures become significantly stronger at 300 and 600 K. For example, a distinct band emerges in the 3.5-6 $\upmu$m region, absent in most harmonic spectra but observed in AIBs (Jourdain de Muizon et al., 1986; Sloan et al., 1997). This feature may originate from the redshift of the C-H stretching band in superhydrogenated PAHs due to anharmonic effects (Mackie et al., 2018a; Yang et al., 2020; Sandford et al., 2013), with additional redshift and intensity enhancement likely driven by temperature effects.

### v2

@@ Abstract
Polycyclic aromatic hydrocarbons (PAHs) are key contributors to interstellar aromatic infrared (IR) bands. However, current spectral databases for IR emission analysis are limited by the omission of vibrational anharmonicity and temperature effects, primarily because of the high computational cost of conventional quantum chemical calculations (QCCs). In this work, we present a machine learning-based molecular dynamics (MLMD) approach that efficiently computes anharmonic IR spectra while incorporating temperature effects. MLMD achieves predictive accuracy comparable to that of QCCs but with significantly reduced computational cost, scaling linearly with the number of atoms in the system. We applied MLMD to calculate the anharmonic spectra of 1 704 PAHs in the NASA Ames PAH IR Spectroscopic Database with up to 216 carbon atoms, demonstrating its capability for high-throughput spectral calculations of large molecular systems. Our results highlight MLMD’s potential to enable the development of extensive molecular spectral datasets, enhancing data-driven analyses of astronomical IR spectra, particularly in anticipation of upcoming data from the James Webb Space Telescope.
@@ I Introduction
In the mid-1970s, distinct infrared (IR) signals were identified in the universe, notably at wavelengths of 3.3, 6.2, 7.7, 8.6, 11.2, and 12.7 $\upmu$m. These signals were later linked to the vibrations of C-H and C-C bonds in aromatic molecules, known as aromatic IR bands (AIB) (Leger & Puget, 1984; Allamandola et al., 1985). Such emissions are primarily attributed to polycyclic aromatic hydrocarbons (PAHs) after the absorption of ultraviolet (UV) photons. PAHs are recognized as key contributors to the evolution of the interstellar medium, influencing processes such as gas heating, ionization balance, and star formation (Tielens, 2008). Determining the population, composition, size distribution, charge state, and chemical structures of interstellar PAHs provides valuable insight into the physical conditions of their host galaxies. However, this remains a significant challenge because of the structural diversity of PAHs. Variations in carbon-ring arrangements, side groups, substitutions, charge states, and isotopologues lead to distinct IR emission characteristics, each governed by specific vibration modes (Peeters et al., 2021).
In order to tackle these challenges, quantum chemical calculations (QCCs) are now an indispensable aid in AIB analysis. Using QCCs, publicly available spectral databases have been established for AIB analysis, such as the NASA Ames PAH IR Spectroscopic Database (PAHdb) (Mattioda et al., 2020; Bauschlicher et al., 2018; Boersma et al., 2014). The availability of these databases enables extensive data-driven analysis of AIBs, providing a more robust interpretation of their features (Boersma et al., 2013, 2015; Maragkoudakis et al., 2020; Sadjadi et al., 2015; Cami, 2011; Shannon & Boersma, 2019; Ricca et al., 2021). Furthermore, these databases have facilitated the analysis of astrochemical components using machine learning (ML) models, demonstrating significant computational efficiency (Kovács et al., 2020) and helping to uncover relationships between molecular structures and spectra (Meng et al., 2021, 2023). Despite these advancements, current theoretical spectral databases face three primary limitations:
@@
  • Neglect of temperature effect: AIB features are typically attributed to IR emission from far-UV-pumped PAHs (Leger & Puget, 1984), where an absorbed UV photon can elevate the effective molecular temperature to several hundred or even thousands of kelvin, an effect that is not taken into account in most theoretical spectral data.
@@
These challenges arise mainly from the high computational cost of QCCs. In particular, the computation time for anharmonic molecular spectra using second-order vibrational perturbation theory (VPT2) is prohibitively long, typically only accommodating molecules with $N_{\text{C}}<25$, where $N_{\text{C}}$ represents the number of carbon atoms in the system (Mackie et al., 2016; Esposito et al., 2024a). The cost of incorporating temperature effects into these calculations using the Wang-Landau random walk algorithm is even more prohibitive, being orders of magnitude higher due to the need for extensive sampling (Chen, 2018; Chen et al., 2019). Facing the vast diversity of PAH species, these limitations significantly hinder the accurate interpretation of observational data, posing substantial challenges in the era of advanced observational facilities such as the James Webb Space Telescope (JWST) (Boersma et al., 2023; Ricca et al., 2024).
@@ II.1 Training Data
The chemical structures of 687 neutral PAH molecules were extracted from the theoretical dataset of PAHdb (v3.2, Mattioda et al. (2020)) to construct the training datasets. The selected PAHs encompass a diverse array of chemical structures that are of astronomical significance. Classical MD simulations were performed using a custom code to generate random configurations of the atoms in each molecule at 300 K, using the adaptive interatomic reactive empirical bond order potential (Stuart et al., 2000). In these simulations, the molecules reached thermal equilibrium in 200 ps using a canonical Nose-Hoover thermostat, with a time step of 0.5 fs. Ten extended structures were extracted from the atomic trajectories of each molecule. As these geometries do not necessarily fall near equilibrium configurations, random Gaussian perturbations were applied to the ground-state atomic coordinates to create 15 more configurations for each molecule to improve the generalizability of the model. Finally, in total, 17 175 configurations (also called molecular conformations) were generated.
@@ II.2 NNFF
NNFF utilizes a deep residual network architecture (ResNet) (He et al., 2016a), implemented within the JAX framework (Bradbury et al., 2018). A ResNet employs skip connections to learn the residual function between the input and output layers, effectively addressing the data degradation problem that arises with increased network depth (He et al., 2016b). This feature is particularly beneficial for modeling complex molecules such as large PAHs, as it enables the capture of deep dependencies and interaction patterns between atoms through multilevel learning, which ultimately leads to more accurate predictions of physicochemical properties (Zhang et al., 2024; Xue et al., 2024).
@@
$\begin{split}L&=0.99\sum^{N_{\text{mol}}}\left\langle\frac{0.2}{N_{\text{atom}}}\sum^{N_{\text{atom}}}\mathrm{ln}\left[\mathrm{cosh}\left(\frac{\left\|\mathbf{f}_{\text{pred}}-\mathbf{f}_{\text{ref}}\right\|_{2}}{0.2}\right)\right]\right\rangle\\ &+0.01\sum^{N_{\text{mol}}}\left\langle 0.02\ \mathrm{ln}\left[\mathrm{cosh}\left(\frac{E_{\text{pred}}-E_{\text{ref}}}{N_{\text{atom}}\times 0.02}\right)\right]\right\rangle,\end{split}$ (1)
where $N_{\text{mol}}$ denotes the total number of molecules, and ${N_{\text{atom}}}$ represents the total number of atoms in a molecule. The subscripts ${}_{\text{pred}}$ and ${}_{\text{ref}}$ denote the values predicted by NNFF and DFT, respectively. $\left\|\mathbf{f}_{\text{pred}}-\mathbf{f}_{\text{ref}}\right\|_{2}$ represents the Euclidean norm of the vector difference between the predicted force and the reference force. The model adopts a core-width sequence of 64:32:16:16 and employs the fully non-linear VeLO optimization technique (Metz et al., 2022) for 600 epochs of training to enhance convergence (Carrete et al., 2023).
To evaluate the performance of the model, we conducted tests on a set of configurations for 3 580 PAHs that were excluded from the training set. The results presented in Figure 2b demonstrate that the NNFF model accurately reproduces the forces calculated by DFT, achieving a coefficient of determination (${R^{2}}$) of approximately 0.99 and a mean root mean square error (RMSE) of 0.00117 hartree/bohr (see Table 1), indicating strong generalizability to unseen samples. This capability allows the NNFF model to effectively replace DFT electronic structure calculations in AIMD, providing robust support for subsequent IR spectrum calculations.
@@ II.3 EPNN
where $p^{\left(\alpha\right)}_{i,\text{pred}}$ and $p^{\left(\alpha\right)}_{i,\text{ref}}$ represent the predicted and reference values of the $\alpha$-th Cartesian component of $\mathbf{p}$ for the $i$-th molecule. To evaluate the performance of the model, we predicted the $\mathbf{p}_{\text{ref}}$ of 3 580 PAH conformations excluded from the training set. As shown in Figure 2c, there is strong agreement between the predicted and reference values, with a ${R^{2}}$ reaching 0.97 and a mean RMSE of 0.0545 debye, as shown in Table 1.
@@ II.4 Anharmonic IR spectrum
To compute the anharmonic IR spectrum of a PAH, MD simulations were performed to simulate its finite-temperature dynamics based on atomic forces predicted by the trained NNFF. We used the MD implementation in the Atomic Simulation Environment (ASE) (Larsen et al., 2017). Initially, the molecule reaches thermal equilibrium in 200 ps in a canonical ensemble (NVT) at a target temperature controlled by the Nose-Hoover thermostat, with a timestep of 0.5 fs. The system was then further simulated in a microcanonical ensemble (NVE) for 200 ps, recording atomic configurations at every 1.0 fs during vibration.
These configurations serve as input to the EPNN, which is used to compute $\mathbf{p}(t)$ at each time step of the vibration. The dipole time autocorrelation function $\langle\mathbf{p}(0)\cdot\mathbf{p}(t)\rangle$, characterizing the time evolution of $\mathbf{p}$, is then subjected to a Fourier transform to obtain the IR intensity(berens1981):
@@
where $\omega$ denotes the angular frequency, $k_{\text{B}}$ represents the Boltzmann constant, $\hbar$ is the reduced Planck constant, $c$ is the speed of light, $T$ indicates the temperature, and $t$ is time.
@@ II.5 Reference methods for comparison
To evaluate the accuracy of our model, we compared it with two established QCC methods to calculate the IR spectra of PAHs. The first method, designated as DFT harmonic, is a widely adopted hybrid DFT approach for harmonic spectrum calculations. In this method, the vibrational frequencies (normal modes) are calculated from the second derivatives of the potential energy with respect to the positions of the nuclei at the stationary point of the system. IR intensities are determined using the double harmonic approximation, which involves computing the derivatives of $\mathbf{p}$ with respect to the normal modes (Langhoff, 1996). The calculations were carried out at the B3LYP/4-31G level of theory using Gaussian 16, ensuring consistency with the QCCs employed in our training dataset. It is important to note that the assumption of a harmonic potential frequently underestimates fundamental frequencies, often necessitating empirical scaling factors to align the computational results with the experimental data (Bauschlicher & Langhoff, 1997). Despite the transferability issues introduced by these empirical scaling factors, this method remains the most commonly utilized method for molecular IR spectrum calculations due to its relatively low computational cost. For the sake of a fair comparison, we do not apply any scaling factors to the DFT harmonic results, as these factors should ideally be derived from fitting to experimental spectra. The DFT harmonic method was used to compute the harmonic spectra for 49 PAH molecules, for which the experimentally measured spectra are available in the PAHdb.
@@
The second method used for comparison is the VPT2 anharmonic spectrum calculations combined with B3LYP/6-311+g* DFT, referred to here as DFT anharmonic. This approach represents the state-of-the-art for computing the molecular anharmonic IR spectra, and has been implemented in Gaussian 16. It requires the computation of the second, third, and fourth derivatives of the potential, which results in significant computational demands (Nielsen, 1951; Mackie et al., 2015). Due to this cost, we applied VPT2 to calculate the anharmonic spectra of 26 PAHs, with the maximum $N_{C}$ being 32, from the 49 PAHs used for the DFT harmonic method.
@@
The 49 PAHs for which the experimental spectra are available in PAHdb (v3.2), ranging from 10 to 50 carbon atoms, exhibit a variety of chemical structures, as detailed in Appendix I. The experimental IR spectra were obtained using matrix isolation techniques at low temperatures (Hudgins & Allamandola, 1995; Hudgins & Sandford, 1998a, b, c; Hudgins & Allamandola, 1999; Mattioda et al., 2003, 2005, 2014, 2017). To facilitate comparison with experimental data, we employed theoretical spectra computed using MLMD at an effective temperature of 50 K, which accounts for molecular vibrations under experimental conditions (Esposito et al., 2024b). In the comparison between computed and experimental spectra, discrete infrared spectral lines were Lorentzian broadened with a full width at half maximum of $18\text{cm}^{-1}$.
@@ III Results and discussion
Figure 3: Comparison of IR spectra for five PAHs computed using the MLMD at 50 K(left, red), DFT harmonic (middle, green), and DFT anharmonic (right, blue) methods, alongside the experimentally measured spectra (black). The intensities are normalized to the maximum values across the four spectra for ease of comparison. The RMSE is computed from the original intensity values, with units of $10^{5}\text{cm}^{2}/\text{mol}$.
@@
This dataset is expected to be a valuable resource for the spectral decomposition of AIBs and for data-driven studies investigating the precise structure-spectral relationships of PAHs. In addition, it will aid in the exploration of temperature effects. For instance, Figure 7 displays the anharmonic IR spectra of a simple mixture of the 1 704 neutral PAHs at 50 K (blue), 300 K (green), and 600 K (red). These spectra represent the sum of individual PAH spectra and are compared to the ground-state harmonic spectra from PAHdb (gray). Notably, the IR signals are significantly enhanced at high temperatures, exhibiting stronger peaks and an elevated plateau. A redshift in peak positions is also observed at high temperatures as a general trend. Furthermore, some emission bands that are weak at lower temperatures become significantly stronger at 300 and 600 K. For example, a distinct band emerges in the 3.5-6 $\upmu$m region, absent in most harmonic spectra but observed in AIBs (Jourdain de Muizon et al., 1986; Sloan et al., 1997). This feature may originate from the redshift of the C-H stretching band in superhydrogenated PAHs due to anharmonic effects (Mackie et al., 2018a; Yang et al., 2020; Sandford et al., 2013), with additional redshift and intensity enhancement likely driven by temperature effects.